OpenStudy (anonymous):

Jack is standing next to a monument in the city square. His shadow is 18 feet long. The shadow of the monument is 150 feet long. If Jack is 6 feet tall, how tall is the monument? 25 feet 50 feet 75 feet 450 feet

OpenStudy (saifoo.khan):

Shadow : Original 18 : 6 150 : x

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (saifoo.khan):

x = 50ft

OpenStudy (katrinakaif):

Use a proportion: 6/x = 18/ 150 18x = 900 x = 50

OpenStudy (anonymous):

x/6 = 150/18 Solve x = 50